# arXiv:2603.00916v1[gr-qc] 01 Mar 2026

### v1

@@ I Introduction
Similarly, some BW models addressed the cosmological constant problem and provided various insights regarding dark matter and energy. For instance, the impacts associated with dark matter could be comprehended using gravity leaking from the extra dimension Heydar-Fard:2007ahl ; Lue:2002sw ; Maartens:2010ar . Furthermore, in the BW context, numerous studies considered the black strings solution rather than a simple extension of the five-dimensional spacetime metric (for more details in terms of BW scenarios like their thermodynamic, geometry, effect of dimensional crossover, and their shapes, check Heydar-Fard:2007ahl ; Lue:2002sw ; Maartens:2010ar ; Chamblin:1999by ; Emparan:1999wa ; Kudoh:2003xz ; Abdolrahimi:2012qi ; Heydarzade:2017xbb ; Shahzad:2021nqt ; Harko:2004ui ; Dadhich:2000am ; Casadio:2003vk ; Hollands:2012sf ; Li:2015vqa ). For instance, the black string solution is more straightforward than the localized five-dimensional BH. It maintains the symmetry of the four-dimensional BH along the extra dimension, making it easier to investigate theoretically. Similarly, one can obtain the black string solution analytically compared to the fully localized five-dimensional BHs, which require numerical techniques.
@@
Duan’s topological current $\phi$-mapping theory is crucial for adopting a topological perspective in thermodynamics. In Refs. 18a ; 19a , it has been suggested that there are two separate approaches for examining topological thermodynamics, focusing on the generalized free energy and temperature function. The temperature approach breaks down the function of the temperature by eradicating the pressure and incorporating auxiliary and topological parameters expressed as $1/\sin\Theta$. These ideas serve as the foundation for the potential’s further construction. In the other technique, the generalized free energy function treats BHs as defects within the thermodynamic parameter space. Their solutions is investigated by utilizing the generalized off-shell free energy, where the instability and stability of BH solutions interpreted by negative and positive winding numbers, respectively (To better understand the thermodynamic topology by utilizing generalized free energy in order to obtain topological charge or winding number, check Refs. 18a ; 19a ; 20a ; 21a ; 23a ; 24a ; 25a ; 26a ; 28a ; 30a ; 31a ; 33a ; 34a ; 35a ; 38a ; 39a ; 40a ; 41a ; 42a ; 43a ; 44a ; 46a ; Bhattacharya:2024bjp ; Yasir:2024wir ; 5000 ; 5001 ; 5002 ; 5003 ; 5004 ; 5005 ; 5006 ).
@@ III Thermal Stability of Brane-World BH through the Barrow entropy
Moreover, in Fig. 2, we have graphically presented the behavior of bound points and divergence concerning the dark matter parameter $\beta$ by using $\delta=0$ (red curve), $\delta=0.3$ (black curve), $\delta=0.7$ (blue curve) and $\delta=1$ [^9] For divergency, we have inserted the $\delta=0.9$ instead of $\delta=1$ in Fig. 2(2(c)) because at $\delta=1$ divergency root becomes infinite. (grey curve). From Fig. 2, it is apparent that stability rises as the dark matter parameter $\beta$ increases and decreases with lower values, indicating that inconsistency increases for the small values of dark matter parameter $\beta$. Furthermore, the divergence shown in Fig. 2(2(c)) is absent in the Bekenstein-Hawking entropy case, demonstrating that it is not an inherent property of the BH. Rather than being generic, the observed behavior arises from the non-additive Barrow entropy modifications, which significantly alter the entropy-temperature correspondence and play a dynamic role when combined with the BW dark-matter parameter $\beta$. Furthermore, these divergences in the case of the Barrow entropy  are verified using thermodynamic geometry. From both Fig. 2, it can be deduced that the zero point (or divergence) of heat capacity $C$ can be avoided for large values of $\alpha$ and $\beta$.
@@ V Topological interpretation of the Brane-world Black holes
For instance, as illustrated in Fig. 4(4(b)), when we set the free parameters to ($\delta=0.3$), ($\alpha=0.7$), and ($\beta=-1.4$), we find two topological charges $(\omega=~+1,~-1)$ with a total topological charge of (W=0W=0) [^11] One can obtain the total topological charge by adding the winding number which is corresponding to the number of phase transition or ZPs 50a . For example, in our case, we have obtained two ZPs or phase transitions corresponding to which we have winding numbers -1 and 1. Thereby, by adding these winding numbers, we can obtain a total topological charge of $W=0$.. However, by reducing the parameter ($\beta$), the number of topological charges also decreases. This is evident in Fig. 4(4(d)), where, with ($\delta$) and ($\alpha$) held constant and ($\beta$) reduced, the number of topological charges decreases to one $(\omega=~-1)$, resulting in a total topological charge of (W=1W=~-1). Interestingly, increasing the parameter ($\delta$), as shown in Fig. 4(4(b)), or altering the parameter ($\alpha$), as depicted in Figs. 5(5(b)) and 5(5(d)), do not affect the number of topological charges. This shows the role of the parameter ($\beta$) in influencing the topological charges, as further illustrated in Figs. 4 and 5. These findings are summarized in Table 2. A similar result is observed when the deformation parameter ($\delta$=0), as shown in Figs. 6(6(b)) and 6(6(c)), highlighting the significant impact of the parameter ($\beta$) on the number of topological charges in the current structure. In Figs. 4(4(a)), 5(4(a)), 6(6(a)), we have plotted the trajectory corresponding to Eq. (33) across various values of the free parameters. When mapping the free energy function as a scalar quantity in the two-dimensional space $(r_{e},\Theta)$, the resulting vector field $(\phi)$ is formulated in such a way that the extremum points of the free energy function act as ZPs for this vector field. In this context, the behavior of the field lines of $(\phi)$ around these ZPs, depending on whether they correspond to a maximum or minimum in the free energy function, exhibits rotational characteristics. The structure of these rotations allows for the assignment of a topological charge to each of these ZPs 19a . Preliminary studies on fundamental BHs, such as Schwarzschild and Reissner-Nordstr’́om BHs, have demonstrated that the total topological charge for these BHs follows a distinct pattern: For a Schwarzschild BH, the total topological charge is W=1W=-1. For a Reissner-Nordström BH, the total topological charge is $W=0$. For an AdS-Reissner-Nordström BH, the total topological charge is $W=+1$ 19a . Since these BHs serve as classical foundations in BH physics, their behavior is utilized as a general model for categorizing BH dynamics, including thermodynamics. As shown in Table (3), BHs with a total topological charge of 1-1 exhibit Schwarzschild-like behavior, whereas those classified with $W=0$ belong to the Reissner-Nordström category. Accordingly, our calculations yield a range of results that, while diverse across different parameter intervals, confirm that within various free parameter domains in the BW scenario, BH behavior aligns well with the established topological charge classifications. Specifically, in our model, W=1W=-1 corresponds to Schwarzschild-like behavior, whereas $W=0$ matches the characteristics of a Reissner-Nordström BH, as demonstrated in Table (2) and in these two cases, considering the shape of the $\tau$ diagram, it seems unlikely that a first-or second-order Van der Waals phase transition will occur in the system.
@@
Figure 4: In Fig. 4, rainbow color arrows presented the normalized vector in $r_{\mathrm{e}}-\Theta$ plane for the Barrow entropy. The trajectory defined by Eq. (33) is showed in Figs. 4(4(a))). In Figs. 4(4(b)), 4(4(c)), 4(4(d)), the ZPs or (phase transition) are demonstrated at coordinates $(r_{e},\Theta)$ along the circular paths which are determined by the free parameter associated with the Barrow entropy.
Figure 5: In Fig. 5, rainbow color arrows presented the normalized vector in $r_{\mathrm{e}}-\Theta$ plane for the Barrow entropy. A trajectory defined by Eq. (33) is showed in Figs. 5(5(a))). While, in Figs. 5(5(b)), 5(5(c)) and 5(5(d)), the ZPs or (phase transition) are demonstrated at coordinates $(r_{e},\Theta)$ along the close paths which are computed by the free parameter associated with the Barrow entropy.
Figure 6: In Fig. 6, rainbow color arrows depicted the normalized vector in $r_{\mathrm{e}}-\Theta$ plane for the Hawking-Bekenstein entropy. The curve defined by Eq. (33) is depicted in Figs. 6(6(a))). In Figs. 6(6(b)) and 6(6(c)), the ZPs (phase transition) are displayed at coordinates $(r_{e},\Theta)$ along the circular paths, which are determined by the free parameter when $\delta=0$.
@@
Kerr BH 41a 0
Kerr-Newman BH 41a 0
$d=5$ Singly Rotating Kerr BH 41a 0
$d\geq 6$ Singly Rotating Kerr BH 41a -1
@@
By evaluating the winding number across parameter space, one can systematically predict the emergence, stability, and classification of black hole phases. The approach not only distinguishes between first- and second-order phase transitions but also generalizes to broader classes of black holes in modified gravity. Ultimately, the topological winding number acts as a robust invariant that encodes the global structure of thermodynamic phase space, offering a powerful diagnostic for understanding black hole thermodynamics in extended theories of gravity. We indeed established that all three primary parameters—the cosmological parameter $\alpha$, the quantum-gravity deformation parameter $\delta$, and the dark matter parameter $\beta$—play crucial and non-trivial roles in shaping the local thermodynamic behavior of the braneworld black hole. As detailed in our study of heat capacity and thermodynamic geometry, each parameter significantly affects the location of stability bounds (roots) and phase transition points (divergences). The Ricci scalar of the Ruppeiner metric, a sensitive probe of microscopic interactions and phase structure, confirmed that singularities shift systematically with variations in $\alpha$, $\delta$, and $\beta$. This demonstrates that the local thermodynamic geometry is a collective function of the entire parameter set. However, when progressing to the global, topological analysis—which classifies the entire thermodynamic phase space via the topological charge WW—a distinct hierarchy of influence emerges. Our systematic investigation reveals the following critical pattern: while varying $\alpha$ or $\delta$ with $\beta$ fixed can alter the characteristics of critical points (e.g., their horizon radius or temperature), these variations do not typically induce a change in the integer value of the total topological charge $W$. The phase space structure, in terms of the number and fundamental type of stable and unstable phases, remains topologically invariant under such changes. Conversely, varying the dark matter parameter $\beta$ proves to be the most potent trigger for a global topological phase transition. For fixed values of $\alpha$ and $\delta$, adjusting $\beta$ beyond critical thresholds reliably switches the total charge between values such as W=1W=-1 and $W=0$. This signifies a fundamental restructuring of the free-energy landscape, where entire solution branches are created or annihilated. The “dominance” of $\beta$ is therefore defined in this specific, rigorous sense: it is the parameter whose variation most directly and consistently governs changes in the global topological invariant of the system, a property more robust and profound than shifting local critical points. This exceptional role of $\beta$ is not a contradiction but a deeper insight stemming from the mathematical structure of the braneworld metric given in Eq. (6). The parameter $\beta$ uniquely appears in a constant, additive term $(-\beta^{2})$. This term acts as a universal shift in the gravitational potential, effectively altering the asymptotic boundary condition for the free energy construction used in the topological method. Parameters like $\alpha$ and $\delta$, while critically important, often enter in a more functional, scale-dependent manner (e.g., $\alpha^{2}r^{2}$). Their variations modify the “shape” of the thermodynamic landscape, but the constant shift from $\beta$ can change its very “foundation,” thereby controlling the topological class. Physically, this underscores that $\beta$, encoding the tidal charge from the extra dimension, impacts the gravitational binding in a more global, integrated manner compared to the more localized or environmental effects of $\alpha$ (cosmological background) and $\delta$ (horizon microstructure). Thereby, our findings present a coherent two-tiered picture: 1. Thermodynamics (Heat Capacity, Ruppeiner Geometry): Governed by a strong, interconnected interplay of all parameters $\alpha$, $\delta$, and $\beta$.
@@ Acknowledgments
The work of KB was supported by the JSPS KAKENHI Grant Numbers JP21K03547, 24KF0100, and Competitive Research Funds for Fukushima University Faculty (25RK011). We would like to sincerely thank Dr. Krishnakanta Bhattacharya for the constructive and insightful discussion.
@@
Figure 8: Fig. 8(8(a)) and 8(8(c)): The allowed range of $\beta$ to have a positive real horizon in BW BH. Fig. 8(8(b)): Metric function with different $\beta$ for BW BH. In Fig. 8(8(b)), we have presented the metric function by putting the dark matter parameter $\beta=0.001$ (green curve), $\beta=0.0025$ (light blue dashed curve), $\beta=0.05$ (dashed green curve), $\beta=0.059$ (dotted dashed orange curve), and $\beta=0.09$ (dotted red curve). In Fig. 8(8(d)), we have plotted the metric function by using $\beta=-1.5$ (green curve), $\beta=-1$ (dotted dashed orange curve), $\beta=0.5$ (dashed red curve), and $\beta=1.5$ (dotted purple curve).

### v2

# Thermodynamic Topology and Photon Spheres Analysis of Black Holes in Brane-World: Insights from Barrow Entropy

@@ I Introduction
Similarly, some BW models addressed the cosmological constant problem and provided various insights regarding dark matter and energy. For instance, the impacts associated with dark matter could be comprehended using gravity leaking from the extra dimension Heydar-Fard:2007ahl ; Lue:2002sw ; Maartens:2010ar . Furthermore, in the BW context, numerous studies considered the black strings solution rather than a simple extension of the five-dimensional spacetime metric (for more details in terms of BW scenarios like their thermodynamic, geometry, effect of dimensional crossover, and their shapes, check Heydar-Fard:2007ahl ; Lue:2002sw ; Maartens:2010ar ; Chamblin:1999by ; Emparan:1999wa ; Kudoh:2003xz ; Abdolrahimi:2012qi ; Heydarzade:2017xbb ; Shahzad:2021nqt ; Harko:2004ui ; Dadhich:2000am ; Casadio:2003vk ; Hollands:2012sf ; Li:2015vqa ; Jawad:2023ypn ). For instance, the black string solution is more straightforward than the localized five-dimensional BH. It maintains the symmetry of the four-dimensional BH along the extra dimension, making it easier to investigate theoretically. Similarly, one can obtain the black string solution analytically compared to the fully localized five-dimensional BHs, which require numerical techniques.
@@
Duan’s topological current $\phi$-mapping theory is crucial for adopting a topological perspective in thermodynamics. In Refs. 18a ; 19a , it has been suggested that there are two separate approaches for examining topological thermodynamics, focusing on the generalized free energy and temperature function. The temperature approach breaks down the function of the temperature by eradicating the pressure and incorporating auxiliary and topological parameters expressed as $1/\sin\Theta$. These ideas serve as the foundation for the potential’s further construction. In the other technique, the generalized free energy function treats BHs as defects within the thermodynamic parameter space. Their solutions is investigated by utilizing the generalized off-shell free energy, where the instability and stability of BH solutions interpreted by negative and positive winding numbers, respectively (To better understand the thermodynamic topology by utilizing generalized free energy in order to obtain topological charge or winding number, check Refs. 18a ; 19a ; 20a ; 21a ; 23a ; 24a ; 25a ; 26a ; 28a ; 30a ; 31a ; 33a ; 34a ; 35a ; 38a ; 39a ; 40a ; 41a ; 42a ; 43a ; 44a ; 46a ; Bhattacharya:2024bjp ; Yasir:2024wir ; 5000 ; 5001 ; 5002 ; 5003 ; 5004 ; 5005 ; 5006 ; Wu:2022whe ; Zhu:2024jhw ; Wu:2024asq ; Chen:2025nto ; Ai:2025vno ; Wu:2023xpq ; Wu:2023fcw ; Chen:2024atr ; Liu:2025iyl ; Chen:2025fse ; Wu:2025wpz ; Rani:2025mip ).
@@ III Thermal Stability of Brane-World BH through the Barrow entropy
Moreover, in Fig. 2, we have graphically presented the behavior of bound points and divergence concerning the dark matter parameter $\beta$ by using $\delta=0$ (red curve), $\delta=0.3$ (black curve), $\delta=0.7$ (blue curve) and $\delta=1$ [^9] For divergency, we have inserted the $\delta=0.9$ instead of $\delta=1$ in Fig. 2(2(c)) because at $\delta=1$ divergency root becomes infinite. (grey curve). From Fig. 2, it is apparent that stability rises as the dark matter parameter $\beta$ increases and decreases with lower values, indicating that inconsistency increases for the small values of dark matter parameter $\beta$. Furthermore, the divergence shown in Fig. 2(2(c)) is absent in the Bekenstein-Hawking entropy case, demonstrating that it is not an inherent property of the BH. Rather than being generic, the observed behavior arises from the non-additive the Barrow entropy  modifications, which significantly alter the entropy-temperature correspondence and play a dynamic role when combined with the BW dark-matter parameter $\beta$. Furthermore, these divergences in the case of the Barrow entropy  are verified using thermodynamic geometry. From both Fig. 2, it can be deduced that the zero point (or divergence) of heat capacity $C$ can be avoided for large values of $\alpha$ and $\beta$.
@@ V Topological interpretation of the Brane-world Black holes
For instance, as illustrated in Fig. 4(4(b)), when we set the free parameters to ($\delta=0.3$), ($\alpha=0.7$), and ($\beta=-1.4$), we find two topological charges $(\omega=~+1,~-1)$ with a total topological charge of (W=0W=0) [^11] One can obtain the total topological charge by adding the winding number which is corresponding to the number of phase transition or ZPs 50a . For example, in our case, we have obtained two ZPs or phase transitions corresponding to which we have winding numbers -1 and 1. Thereby, by adding these winding numbers, we can obtain a total topological charge of $W=0$.. However, by reducing the parameter ($\beta$), the number of topological charges also decreases. This is evident in Fig. 4(4(d)), where, with ($\delta$) and ($\alpha$) held constant and ($\beta$) reduced, the number of topological charges decreases to one $(\omega=~-1)$, resulting in a total topological charge of (W=1W=~-1). Interestingly, increasing the parameter ($\delta$), as shown in Fig. 4(4(b)), or altering the parameter ($\alpha$), as depicted in Figs. 5(5(b)) and 5(5(d)), do not affect the number of topological charges. This shows the role of the parameter ($\beta$) in influencing the topological charges, as further illustrated in Figs. 4 and 5. These findings are summarized in Table 2. A similar result is observed when the deformation parameter ($\delta$=0), as shown in Figs. 6(6(b)) and 6(6(c)), highlighting the significant impact of the parameter ($\beta$) on the number of topological charges in the current structure. In Figs. 4(4(a)), 5(4(a)), 6(6(a)), we have plotted the trajectory corresponding to Eq. (33) across various values of the free parameters. When mapping the free energy function as a scalar quantity in the two-dimensional space $(r_{e},\Theta)$, the resulting vector field $(\phi)$ is formulated in such a way that the extremum points of the free energy function act as ZPs for this vector field. In this context, the behavior of the field lines of $(\phi)$ around these ZPs, depending on whether they correspond to a maximum or minimum in the free energy function, exhibits rotational characteristics. The structure of these rotations allows for the assignment of a topological charge to each of these ZPs 19a . Preliminary studies on fundamental BHs, such as Schwarzschild and Reissner-Nordström BHs, have demonstrated that the total topological charge for these BHs follows a distinct pattern: For a Schwarzschild BH, the total topological charge is W=1W=-1. For a Reissner-Nordström BH, the total topological charge is $W=0$. For an AdS-Reissner-Nordström BH, the total topological charge is $W=+1$ 19a . Since these BHs serve as classical foundations in BH physics, their behavior is utilized as a general model for categorizing BH dynamics, including thermodynamics. As shown in Table (3), BHs with a total topological charge of 1-1 exhibit Schwarzschild-like behavior, whereas those classified with $W=0$ belong to the Reissner-Nordström category. Accordingly, our calculations yield a range of results that, while diverse across different parameter intervals, confirm that within various free parameter domains in the BW scenario, BH behavior aligns well with the established topological charge classifications. Specifically, in our model, W=1W=-1 corresponds to Schwarzschild-like behavior, whereas $W=0$ matches the characteristics of a Reissner-Nordström BH, as demonstrated in Table (2) and in these two cases, considering the shape of the $\tau$ diagram, it seems unlikely that a first-or second-order Van der Waals phase transition will occur in the system.
@@
Figure 4: In Fig. 4, rainbow color arrows presented the normalized vector in $r_{\mathrm{e}}-\Theta$ plane for the Barrow entropy. The trajectory defined by Eq. (33) is showed in Fig. 4(4(a)). In Figs. 4(4(b)), 4(4(c)), 4(4(d)), the ZPs or (phase transition) are demonstrated at coordinates $(r_{e},\Theta)$ along the circular paths which are determined by the free parameter associated with the Barrow entropy.
Figure 5: In Fig. 5, rainbow color arrows presented the normalized vector in $r_{\mathrm{e}}-\Theta$ plane for the Barrow entropy. A trajectory defined by Eq. (33) is showed in Fig. 5(5(a)). While, in Figs. 5(5(b)), 5(5(c)) and 5(5(d)), the ZPs or (phase transition) are demonstrated at coordinates $(r_{e},\Theta)$ along the close paths which are computed by the free parameter associated with the Barrow entropy.
Figure 6: In Fig. 6, rainbow color arrows depicted the normalized vector in $r_{\mathrm{e}}-\Theta$ plane for the Hawking-Bekenstein entropy. The curve defined by Eq. (33) is depicted in Fig. 6(6(a)). In Figs. 6(6(b)) and 6(6(c)), the ZPs (phase transition) are displayed at coordinates $(r_{e},\Theta)$ along the circular paths, which are determined by the free parameter when $\delta=0$.
@@
Kerr BH Wu:2022whe ; 41a 0
Kerr-Newman BH Wu:2022whe ; 41a 0
$d=5$ Singly Rotating Kerr BH Wu:2022whe ; 41a 0
$d\geq 6$ Singly Rotating Kerr BH Wu:2022whe ; 41a -1
@@
By evaluating the winding number across parameter space, one can systematically predict the emergence, stability, and classification of black hole phases. The approach not only distinguishes between first- and second-order phase transitions but also generalizes to broader classes of black holes in modified gravity. Ultimately, the topological winding number acts as a robust invariant that encodes the global structure of thermodynamic phase space, offering a powerful diagnostic for understanding black hole thermodynamics in extended theories of gravity. We indeed established that all three primary parameters—the cosmological parameter $\alpha$, the quantum-gravity deformation parameter $\delta$, and the dark matter parameter $\beta$—play crucial and non-trivial roles in shaping the local thermodynamic behavior of the BW black hole. As detailed in our study of heat capacity and thermodynamic geometry, each parameter significantly affects the location of stability bounds (roots) and phase transition points (divergences). The Ricci scalar of the Ruppeiner metric, a sensitive probe of microscopic interactions and phase structure, confirmed that singularities shift systematically with variations in $\alpha$, $\delta$, and $\beta$. This demonstrates that the local thermodynamic geometry is a collective function of the entire parameter set. However, when progressing to the global, topological analysis—which classifies the entire thermodynamic phase space via the topological charge WW—a distinct hierarchy of influence emerges. Our systematic investigation reveals the following critical pattern: while varying $\alpha$ or $\delta$ with $\beta$ fixed can alter the characteristics of critical points (e.g., their horizon radius or temperature), these variations do not typically induce a change in the integer value of the total topological charge $W$. The phase space structure, in terms of the number and fundamental type of stable and unstable phases, remains topologically invariant under such changes. Conversely, varying the dark matter parameter $\beta$ proves to be the most potent trigger for a global topological phase transition. For fixed values of $\alpha$ and $\delta$, adjusting $\beta$ beyond critical thresholds reliably switches the total charge between values such as W=1W=-1 and $W=0$. This signifies a fundamental restructuring of the free-energy landscape, where entire solution branches are created or annihilated. The “dominance” of $\beta$ is therefore defined in this specific, rigorous sense: it is the parameter whose variation most directly and consistently governs changes in the global topological invariant of the system, a property more robust and profound than shifting local critical points. This exceptional role of $\beta$ is not a contradiction but a deeper insight stemming from the mathematical structure of the braneworld metric given in Eq. (6). The parameter $\beta$ uniquely appears in a constant, additive term $(-\beta^{2})$. This term acts as a universal shift in the gravitational potential, effectively altering the asymptotic boundary condition for the free energy construction used in the topological method. Parameters like $\alpha$ and $\delta$, while critically important, often enter in a more functional, scale-dependent manner (e.g., $\alpha^{2}r^{2}$). Their variations modify the “shape” of the thermodynamic landscape, but the constant shift from $\beta$ can change its very “foundation,” thereby controlling the topological class. Physically, this underscores that $\beta$, encoding the tidal charge from the extra dimension, impacts the gravitational binding in a more global, integrated manner compared to the more localized or environmental effects of $\alpha$ (cosmological background) and $\delta$ (horizon microstructure). Thereby, our findings present a coherent two-tiered picture: 1. Thermodynamics (Heat Capacity, Ruppeiner Geometry): Governed by a strong, interconnected interplay of all parameters $\alpha$, $\delta$, and $\beta$.
@@ Acknowledgments
The work of KB was supported by the JSPS KAKENHI Grant Numbers JP21K03547, 24KF0100, 25KF0176, and Competitive Research Funds for Fukushima University Faculty (25RK011). We would like to sincerely thank Dr. Krishnakanta Bhattacharya for the constructive and insightful discussion.
@@
Figure 8: Figs. 8(8(a)) and 8(8(c)): The allowed range of $\beta$ to have a positive real horizon in BW BH. Figs. 8(8(b)) and 8(8(d)): Metric function with different $\beta$ for BW BH. In Fig. 8(8(b)), we have presented the metric function by putting the dark matter parameter $\beta=0.001$ (green curve), $\beta=0.0025$ (light blue dashed curve), $\beta=0.05$ (dashed green curve), $\beta=0.059$ (dotted dashed orange curve), and $\beta=0.09$ (dotted red curve). In Fig. 8(8(d)), we have plotted the metric function by using $\beta=-1.5$ (green curve), $\beta=-1$ (dotted dashed orange curve), $\beta=0.5$ (dashed red curve), and $\beta=1.5$ (dotted purple curve).